|IODP Proceedings Volume contents Search|
|Expedition reports Research results Supplementary material Drilling maps Expedition bibliography|
We transited 8.7 h from Site U1328 in dynamic positioning (DP) mode to Site U1329 (proposed Site CAS-05D; Collett et al., 2005) while we downloaded the LWD/MWD data collected at the three previous sites. Hole U1329A was spudded at 1830 h on 25 September 2005, at an estimated water depth of 959.1 meters below sea level (mbsl; 970 meters below rig floor [mbrf]) (see mudline core estimate in "Hole U1329C"). LWD/MWD drilling progressed without incident and within the prescribed protocol guidelines (see "Downhole logging" in the "Methods" chapter). A total depth (TD) of 220 mbsf was reached at 0800 h on 26 September. After displacing the hole with 10.5 pounds per gallon (ppg) mud, the last of the tools and the bit arrived on deck at 1330 h, officially ending operations in Hole U1329A. We offset 15 m to the north to start coring operations in Hole U1329B (Fig. F6).
On the morning of 26 September 2005, we learned that the LWD tool and personnel transfer boat had to return to Coos Bay, Oregon, twice: once to drop off a sick passenger and again because of engine problems. Ultimately, the engine problems were not serious, the transfer boat was quickly repaired, and it left Coos Bay at 1025 h. The exchange boat arrived at 1630 h on 27 September. Tool offloading and personnel transfer were completed, and the boat departed at 2125 h for Aberdeen, Washington. While waiting for the boat and during transfer operations, the drill string was tripped to near the seafloor. Hole U1329B was spudded, with the first core arriving on deck at 2105 h with 10.02 m of recovery, therefore not recovering the mudline. Recovery of the mudline core is important for high-resolution, shallow geochemical and microbiological studies. We terminated Hole U1329B to start a new hole so that we could obtain a mudline core.
The bit was raised 5 m higher than for the Hole U1329B spud position, and without offsetting, Hole U1329C was spudded with core on the deck at 2150 h on 27 September 2005. Core 311-U1329C-1H recovered 8.19 m, providing an estimated seafloor depth of 946.4 mbsl (957.4 mbrf, which indicated that the seafloor depth had been overestimated by >6 m in Holes U1329A and U1392B (Table T1). Coring with the advanced piston corer (APC) system continued to 140.2 mbsf (106% recovery), where we switched to the XCB coring system and advanced the hole to 188.5 mbsf (82% recovery) (Table T1). Five APC temperature measurements were made with the advanced piston corer temperature (APCT) or third-generation advanced piston corer temperature (APCT-3) tools (Table T1). In addition, the APCT-3 housing was run without internal electronics to test water tightness for Core 311-U1329C-2H. Two PCS cores (311-U1329C-7P and 14P) were collected at 55.6 and 114.6 mbsf, respectively. Core 311-U1329C-14P was recovered without pressure. The third PCS core deployment (Core 311-U1329C-23P) returned to the rig floor at 0210 h on 29 September without the cutting shoe, which was left in the hole. Attempts to drill past it were unsuccessful. Coring in Hole U1329C was terminated at 189.5 mbsf; ~30 m shy of the target depth.
Swell from a nearby low pressure system (>5 m heave) made conditions too rough to attempt wireline logging or recover quality piston cores from a new hole; therefore, we decided to wait on the weather to improve before resuming operations. After displacing the hole with 10.5 ppg mud, the drill pipe was pulled to 81.5 mbsf at 0630 h on 29 September. After receiving the forecast for continued poor conditions throughout that day, we decided to pull out of Hole U1329C and drill a dedicated logging hole when better conditions would permit. This scenario had the added benefit of penetration to total depth and a better chance for good hole conditions for logging.
After pulling out of Hole U1329C, we offset 15 m south of Hole U1329A and spudded Hole U1329D at 1115 h on 29 September 2005. At 1955 h, after drilling Hole U1329D to 165.6 mbsf, increased ship heave (>6 m) forced us to suspend operations again. By 2130 h, improved conditions allowed us to resume drilling. After relatively slow drilling (~5 m/h) in the lower ~35 m, we took a single XCB core (311-U1329D-1X; 1.03 m of recovery) at the bottom of Hole U1329D, which was completed to a TD of 210.5 mbsf at 1015 h on 30 September. After reaching TD in Hole U1329D, we switched to the wireline logging program. The first tool string deployed was the standard triple combo, which reached a depth of 209 mbsf. The hole was logged without incident. The second logging run consisted of two passes of the FMS-sonic tool string. On the first run, the FMS-sonic tool string reached a depth of 209 mbsf, but on the second run the FMS-sonic tool string could only reach a depth of 171 mbsf. The caliper log showed that parts of the hole were severely enlarged. After the wireline tools were recovered, the hole was displaced with 10.5 ppg mud, and the pipe was pulled out of the hole, ending operations in Hole U1329D at 0020 h, 1 October.
After offsetting 15 m to the north of Hole U1329C in DP mode, Hole U1329E was spudded at 0235 h on 1 October 2005. Core 311-U1329E-1H recovered 5.05 m, which indicated a seafloor depth of 945.5 mbsl (956.5 mbrf). The hole was advanced with the APC system to 33.5 mbsf, where the PCS was deployed, and Core 311-U1329E-5P was recovered without pressure. After washing down to 54.5 mbsf, Core 311-U1329E-6H was taken to obtain a temperature measurement with the APCT tool. The Davis-Villinger Temperature Probe (DVTP) was deployed immediately after Core 311-U1329E-6H to provide a calibration point with the APCT. The rest of the hole was devoted to pressure coring runs with intervening washed intervals (Table T1). The PCS was deployed two more times at 73.5 and 125.0 mbsf (Cores 311-U1329E-7P and 10P, respectively), which were successfully recovered under pressure. The FPC was deployed at 104.0 mbsf but did not return under pressure. Finally, the HRC was deployed at 114.5 mbsf and retrieved a full core under pressure. After displacing the hole with 10.5 ppg mud, the bottom-hole assembly (BHA) was pulled clear of the seafloor at 2232 h on 1 October, ending operations in Hole U1329E. The drill string was tripped to the rig floor in preparation for the short 9.2 nmi transit to Site U1327.